# Low-Acceleration Gravitational Anomaly from Bayesian 3D Modeling of Wide Binary Orbits: Methodology and Results with Gaia DR3

### v1

@@ 3.1 Overview and inspection of the individual Bayesian results
To sum up, the Bayesian outputs of the parameters $e$, $\phi_{0}$, $i$ and $\log_{10}f_{M}$ are all reasonable, individually and statistically. As for the gravity parameter $\Gamma$, only eight wide binaries listed in Table 2 return unacceptable PDFs given any theoretically conceivable range. These are interpreted as exceptional systems that hide kinematic contaminants. These systems will not be used to derive consolidated values of $\Gamma$ in the main analyses.
@@ 3.2 Consolidated gravity in a Newtonian regime
Possible systematics can be explored by varying the sample selection criteria or the prior inputs. The difference between the thermal prior and the flat prior is moderate, but one can still ask whether a non-thermal prior with $\alpha\neq 1$ in $p(e)=(1+\alpha)e^{\alpha}$ can make a difference. For example, Hwang et al. (2022) suggest $\alpha\approx 1.3$ for wide binaries in a MOND regime. With such a superthermal distribution, the median of $e$ is $\approx 0.74$ differing only by $0.03$ from the thermal case, which is certainly negligible compared with the difference between the thermal and the flat.
@@
Varying the limit of $\sigma_{v_{r}}$ has no impact on the results as the numbers of wide binaries in the nominal and alternative samples are already sufficiently large and all $\sigma_{v_{r}}$ values in the statistical sample are small enough compared with relatively large internal relative velocities in the pairs. Varying the limit of ruwe from the nominal limit $1.3$ has no impact either. Figure 20 shows the results with ruwe $<1.2$ or $<1.4$. Thus, it is concluded that the measured value is $\Gamma=0.000\pm 0.011$ in the Newtonian regime based on the nominal sample and any systematics are negligible.
@@ 3.3 Consolidated gravity in transition and MOND regimes
Second, the deviation is larger in the MOND regime than the transition regime, indicating a gradual deviation from Newton in the high-acceleration regime ($\gtrsim 100a_{0}$) to modified gravity in the low-acceleration regime ($\lesssim a_{0}$). This trend and the estimated value $\Gamma_{\rm{corrected}}=0.085\pm 0.040$ (after correcting for the bias described below) in the MOND regime based on the nominal sample agree well with the generic prediction $\approx 0.07$ of MOND Lagrangian gravity models (e.g. Bekenstein & Milgrom 1984; Milgrom 2010). The value $\Gamma_{\rm{corrected}}=0.121\pm 0.049$ based on the alternative sample is also consistent with $\approx 0.07$. These results through the new Bayesian methodology match well several recent results (Chae, 2023a, 2024a, 2024c; Hernandez, 2023; Hernandez et al., 2024a) based on various statistical analyses utilizing $v_{p}$ only.
@@
The above results are based on the two samples with ruwe $<1.30$. More results with alternative cuts of ruwe can be found in Figure 22. Clearly, for the binaries that have already passed a number of criteria, the ruwe cut does not seem matter up to $2.0$.
@@
Now I describe the systematic bias in $\Gamma$ induced by the RV uncertainties. To estimate it, I use mock Newtonian binaries produced with the procedure described in Section 2.4. For each Gaia binary of the nominal sample ($N_{\rm{binary}}=111$) in the range $2.860.15<s/r_{\rm{M}}<$ 2.86, I produce 100 random binaries with the same values of $s$ and $M_{\rm{tot}}$ (and thus the same $s/r_{\rm{M}}$ as well), and with the same endowed uncertainties of RVs but random otherwise. PDFs of $\Gamma$ are derived for these mock binaries. Thus, a total of $100\times 111$ different mock PDFs of $\Gamma$ are in hand. Now, $N$ sets of mock PDFs of $\Gamma$ corresponding to the set of 111 Gaia binaries are obtained using the mock PDFs in a bootstrap manner. Here, each set is composed of 111 PDFs, each of which is drawn from the set of the ready-made 100 mock PDFs for each Gaia binary. I consider $N=10^{5}$, meaning that the 100 mock PDFs for each Gaia binary are used multiple times. Although any two sets can share some PDFs in doing so, all sets are different and can be taken as random Newtonian realizations.
@@
Figure 23: $N=10^{5}$ mock Newtonian samples corresponding to each Gaia binary sample are used to obtain $N$ consolidated distributions of $\Gamma$. In practice, a bootstrap strategy is employed to save the computing time; only a pool of 100 mock Newtonian binaries are produced for each Gaia binary and these $N_{\rm{binary}}$ pools are used to provide different $N$ samples in a bootstrap way (see the text for the details). Each histogram shows the distribution of the medians ($\Gamma_{\rm{med}}$) of the $N$ consolidated distributions of $\Gamma$. The left column considers three samples in the MOND regime with different upper limits of $\sigma_{v_{r}}$. The samples with $\sigma_{v_{r}}<380$ m s-1 and $<330$ m s-1 are the nominal and alternative samples shown in Figure 21. The smaller sample with $\sigma_{v_{r}}<280$ m s-1 is considered only here for the purpose of illustration. As the upper limit of $\sigma_{v_{r}}$ increases from top to bottom, the histogram exhibits two systematic variations: (1) the median represented by the vertical blue line is gradually shifted to the right from top to bottom panel due to the gradually increased scatters of $v_{r}$, (2) the width of the histogram decreases gradually as the sample size increases. The value $\Gamma_{\rm{med,obs}}$ for each Gaia sample (derived in Figure 21) is indicated by the vertical red line. $P(\Gamma_{\rm{med}}>\Gamma_{\rm{med,obs}})$ is the Newton-predicted probability that the observed value or higher would occur by chance. The right column shows the results for the Gaia samples in the transition+MOND regime. The very small values of $P(\Gamma_{\rm{med}}>\Gamma_{\rm{med,obs}})$ for the nominal and alternative samples show the difficulty of Newtonian gravity. Note, however, that the widths of the histograms in these Newtonian simulations are somewhat different from the widths of the corresponding consolidated distributions for the Gaia samples shown in Figure 21 due to difference in gravity.
@@
For the MOND regime, the Newton-predicted distributions are shifted to the right, and the systematic shift increases (although the width of the distribution gets narrower) as the upper limit of $\sigma_{v_{r}}$ increases. For the MOND+transition regime, the distributions seem to be dominated by more numerous and narrower PDFs in the transition regime, so that the shifts are very small. Each distribution represents a predicted probability distribution of mock observations in a Newtonian world. That is, if gravity were Newtonian, an observed value of $\Gamma_{\rm{med}}$ would fall within a likely range. However, what happens is that the observed values fall outside the predicted likely ranges and thus seem very unlikely in Newtonian gravity as the calculated survival probabilities are indicated in the figure. Here I note that the widths of the simulated distributions are somewhat different from those of the corresponding observed consolidated distributions shown in Figure 21. This is because the simulations were carried out in gravity different from the gravity that the data support.
@@ 4.1 Meaning of the results and an exploration of possible systematic errors
When the center-of-mass of a binary system orbits in the Galaxy, the relative motion between the two stars will suffer from perturbations due to long-range chance encounters with other stars and the Galactic tidal field (see Jaing & Tremaine 2010). Such perturbations can accumulate over a long period of time since the birth of the binary system (e.g. several giga years), leading to an evolved statistical property of binary populations as studied by Jaing & Tremaine (2010) in Newtonian gravity. The maximum projected separation for the wide binaries in the nominal sample (Table 1) is $\approx 25$ kau. This corresponds to $\approx 0.074r_{J}$ where $r_{J}\approx 1.65$ pc is the Jacobi radius for the median total mass of $1.82M_{\odot}$ based on Equation (43) of Jaing & Tremaine (2010). According to Newtonian simulation results shown in Figure 7 of Jaing & Tremaine (2010), presently observed binaries with $s/r_{J}\lesssim 0.074$ would be expected to follow Keplerian behavior well. This indicates that the assumption of elliptical orbits is valid as long as Newtonian gravity is tested with the wide binaries of this work. Thus, the disagreement (Figure 21) with Newton in low-acceleration regimes (along with the agreement in a strong-acceleration regime) is an unacceptable result in Newtonian gravity.
@@
However, in MOND or other modified gravity it is unclear how good the two assumptions will be and how to interpret the Bayesian inferred values of $\Gamma$ ($>0$) (i.e.  $G_{\rm{eff}}>G_{\rm{N}}$). In MOND gravity (e.g. Bekenstein & Milgrom 1984; Milgrom 2010), the rather strong external field $\approx 1.9a_{0}$ of the Galaxy can warp the orbit and the orbit may not be closed when the external field and the rotation axis is inclined (Brada & Milgrom, 2000). In such gravity models, numerical simulations are needed to produce mock wide binaries and to infer $\Gamma$ in the instantaneous motions of the binaries. The values of $\Gamma$ can then be compared with those for the Gaia binaries.
@@
Therefore, a specific modified gravity model cannot be accurately tested with the measured value of $\Gamma$. The agreement of the Bayesian-inferred values of $\Gamma=0.085\pm 0.040$ (MOND regime) and $\Gamma=0.063\pm 0.015$ (transition+MOND regime) with the generic prediction of MOND gravity ($\approx 0.07$) needs to be taken with a grain of salt. In this respect, it would be interesting to reconsider the wide binaries with wildly large values of $\Gamma$ (Table 2). These systems were rejected as systems with suspicious kinematic contaminants because they appeared to be incompatible not only with Newtonian gravity but also with MOND gravity. However, it is unclear whether MOND gravity could accommodate these systems until realistic simulations are used to probe all possible orbital motions. Whatever is the case with MOND gravity, if at least some of these systems were pure binaries without kinematic contaminants, Newtonian gravity would be ruled out more severely.
@@ 4.2 Wide binary results in the context of recent small-scale gravity tests
At the heart of the debate between standard gravity (with dark matter) and MOND or modified gravity (without dark matter) in the universe is whether a direct gravitational boost occurs at low acceleration $\lesssim a_{0}$. The low-acceleration gravitational boost is the fundamental nature that is shared by all models belonging to the nonstandard paradigm, while solar system tests discussed above are more about specifics of models. This means that even if solar system tests truly rule out certain modified gravity models, they cannot rule out the MOND or modified gravity paradigm as long as the low-acceleration gravitational anomaly is experimentally true. This fundamental nature cannot be so easily tested with galactic dynamics (although feasible in principle and will be done eventually) because the prediction of modified gravity needs to be distinguished from that of the cosmologically inferred dark matter density (Chae, 2022). Unlike galactic dynamics, wide binary internal dynamics does not care whether dark matter density is zero or what general relativity infers cosmologically, and thus most directly tests the low-acceleration gravitational anomaly. What the results from this work and several recent studies (Chae, 2023a, 2024a, 2024c; Hernandez, 2023; Hernandez et al., 2024a) show clearly is that the fundamental low-acceleration gravitational boost is present in the currently available data.

### v3

@@ 3.1 Overview and inspection of the individual Bayesian results
To sum up, the Bayesian outputs of the parameters $e$, $\phi_{0}$, $i$ and $\log_{10}f_{M}$ are all reasonable, individually and statistically. As for the gravity parameter $\Gamma$, only eight wide binaries listed in Table 2 return unacceptable PDFs given any theoretically conceivable range. These are interpreted as exceptional systems that hide kinematic contaminants and/or abnormal RV data. These systems will not be used to derive consolidated values of $\Gamma$ in the main analyses. Alternative interpretations are discussed in Section 4.1.
@@ 3.2 Consolidated gravity in a Newtonian regime
Possible systematics can be explored by varying the sample selection criteria or the prior inputs. First of all, the results with the H2022 prior on eccentricities are essentially the same as those with the thermal prior because $\alpha$ in $p(e)=(1+\alpha)e^{\alpha}$ is close to 1 for the Newtonian regime under consideration. Varying the limit of $\sigma_{v_{r}}$ has no impact on the results as the numbers of wide binaries in the nominal and alternative samples are already sufficiently large and all $\sigma_{v_{r}}$ values in the statistical sample are small enough compared with relatively large internal relative velocities in the pairs. Varying the limit of ruwe from the nominal limit $1.3$ has no impact either. Figure 21 shows the results with ruwe $<1.2$ or $<1.4$. Thus, it is concluded that the measured value is $\Gamma=0.000\pm 0.011$ or $\gamma_{g}=1.00\pm 0.05$ in the Newtonian regime based on the nominal sample and any systematics are negligible.
@@ 3.3 Consolidated gravity in transition and MOND regimes
Second, the deviation is larger in the MOND regime than the transition regime, indicating a gradual deviation from Newton in the high-acceleration regime ($\gtrsim 100a_{0}$) to modified gravity in the low-acceleration regime ($\lesssim a_{0}$). This trend and the estimated value $\Gamma_{\rm{corrected}}=0.085\pm 0.040$ or $\gamma_{g}=1.48_{-0.23}^{+0.30}$ (after correcting for the bias described below) in the MOND regime based on the nominal sample agree well with the generic prediction $\Gamma\approx 0.07$ (or $\gamma_{g}\approx 1.4$) of MOND Lagrangian gravity models (e.g. Bekenstein & Milgrom 1984; Milgrom 2010). The value $\Gamma_{\rm{corrected}}=0.121\pm 0.049$ or $\gamma_{g}=1.75_{-0.32}^{+0.50}$ based on the alternative sample is also consistent with $\Gamma\approx 0.07$. These results through the new Bayesian methodology match well several recent results (Chae, 2023a, 2024a, 2024c; Hernandez, 2023; Hernandez et al., 2024a) based on various statistical analyses utilizing $v_{p}$ only.
@@
The above results are based on the two samples with ruwe $<1.30$. More results with alternative cuts of ruwe can be found in Figure 23. Clearly, for the binaries that have already passed a number of criteria, the ruwe cut does not seem to matter up to $2.0$.
Now I describe the systematic bias in $\Gamma$ induced by the RV uncertainties. To estimate it, I use mock Newtonian binaries produced with the procedure described in Section 2.4. The reason for this is twofold. First, to estimate the bias in $\Gamma$, the unbiased value must be accurately known. While the unbiased value is exact (that is, zero) in Newtonian gravity, it is not so in any modified gravity. Second, based on all previous studies of wide binary gravity tests, any effective gravity experienced by wide binaries is not expected to be drastically different from Newton but rather pseudo-Newtonian or something similar. Thus, at least as a first-order approximation, Newtonian gravity can be used for the purpose of estimating the bias due to large uncertainties of RVs. However, I will also consider mock binaries produced with the observed Gaia binaries themselves by giving scatters to the observed RVs. It will be shown that the biases estimated with mock Newtonian and mock Gaia samples are similar. The nominal estimates will be based on mock Newtonian samples.
@@
For each Gaia binary of the nominal sample ($N_{\rm{binary}}=111$) in the range $2.860.15<s/r_{\rm{M}}<$ 2.86, I produce 100 random binaries with the same values of $s$ and $M_{\rm{tot}}$ (and thus the same $s/r_{\rm{M}}$ as well), and with the same endowed uncertainties of RVs but random otherwise. PDFs of $\Gamma$ are derived for these mock binaries. Thus, a total of $100\times 111$ different mock PDFs of $\Gamma$ are in hand. Now, $N$ sets of mock PDFs of $\Gamma$ corresponding to the set of 111 Gaia binaries are obtained using the mock PDFs in a bootstrap manner. Here, each set is composed of 111 PDFs, each of which is drawn from the set of the ready-made 100 mock PDFs for each Gaia binary. I consider $N=10^{5}$, meaning that the 100 mock PDFs for each Gaia binary are used multiple times. Although any two sets can share some PDFs in doing so, all sets are different and can be taken as random Newtonian realizations.
@@
Figure 24: $N=10^{5}$ mock Newtonian samples corresponding to each Gaia binary sample are used to obtain $N$ consolidated distributions of $\Gamma$. In practice, a bootstrap strategy is employed to save the computing time; only a pool of 100 mock Newtonian binaries are produced for each Gaia binary and these $N_{\rm{binary}}$ pools are used to provide different $N$ samples in a bootstrap way (see the text for the details). Each histogram shows the distribution of the medians ($\Gamma_{\rm{med}}$) of the $N$ consolidated distributions of $\Gamma$. The left column considers three samples in the MOND regime with different upper limits of $\sigma_{v_{r}}$. The samples with $\sigma_{v_{r}}<380$ m s-1 and $<330$ m s-1 are the nominal and alternative samples shown in Figure 22. The smaller sample with $\sigma_{v_{r}}<280$ m s-1 is considered only here for the purpose of illustration. As the upper limit of $\sigma_{v_{r}}$ increases from top to bottom, the histogram exhibits two systematic variations: (1) the median represented by the vertical blue line is gradually shifted to the right from top to bottom panel due to the gradually increased scatters of $v_{r}$, (2) the width of the histogram decreases gradually as the sample size increases. The value $\Gamma_{\rm{med,obs}}$ for each Gaia sample (derived in Figure 22) is indicated by the vertical red line. $P(\Gamma_{\rm{med}}>\Gamma_{\rm{med,obs}})$ is the Newton-predicted probability that the observed value or higher would occur by chance. The right column shows the results for the Gaia samples in the transition+MOND regime. The very small values of $P(\Gamma_{\rm{med}}>\Gamma_{\rm{med,obs}})$ for the nominal and alternative samples show the difficulty of Newtonian gravity. Note, however, that the widths of the histograms in these Newtonian simulations are somewhat different from the widths of the corresponding consolidated distributions for the Gaia samples shown in Figure 22 (See the text for further details.).
@@
For the MOND regime, the Newton-predicted distributions are shifted to the right, and the systematic shift increases (although the width of the distribution gets narrower) as the upper limit of $\sigma_{v_{r}}$ increases. For the MOND+transition regime, the distributions seem to be dominated by more numerous and narrower PDFs in the transition regime, so that the shifts are very small. Each distribution represents a predicted probability distribution of mock observations in a Newtonian world. That is, if gravity were Newtonian, an observed value of $\Gamma_{\rm{med}}$ would fall within a likely range. However, what happens is that the observed values fall outside the predicted likely ranges and thus seem very unlikely in Newtonian gravity as the calculated survival probabilities are indicated in the figure. Here, I note that the widths of the simulated distributions agree up to 20% with those of the corresponding observed consolidated distributions shown in Figure 22. The widths of the simulated distributions are somewhat narrower in the MOND regime but somewhat wider in the transition+MOND regime.
Figure 25: Similar to Figure 24 but for $N=10^{5}$ mock “Gaia” samples that are produced by adding to the Gaia RV values random scatters using their reported uncertainties. Thus, in this case the benchmark gravity is not Newton but the Gaia-implied one (uncorrected for the bias being investigated) that is of course uncertain. Unlike Figure 24, the transition regime is exhibited in the left column since there occur small shifts. Because the benchmark gravity models indicated by red vertical lines are subject to uncertainties, the shifts are also unstable as opposed to the Newtonian case in Figure 24 where the shifts are gradual from top to bottom. The widths of the distributions are similar to, or somewhat different from, the corresponding ones in Figure 22 and Figure 24. See the text for further details.
Simulations using Gaia RVs themselves are also carried out in a very similar way. Here the Gaia reported RVs are used to produce mock RVs by adding scatters using the reported nominal uncertainties. Thus, in this case the uncorrected gravity based on the Gaia data provides the input gravity as opposed to Newtonian gravity in the above case. Because the input gravity is not known precisely in this case, there is a minor difference in the details of the procedure. That is to say, because the Gaia-reported RVs are already scattered quantities due to measurement errors, adding additional scatters can, in few cases, produce too wildly scattered RVs that lead to extremely unlikely PDFs of $\Gamma$, e.g. more than $4\sigma$ away from the “input” gravity (note that there was no such concern in the case of Newtonian simulation because RVs exactly follow the underlying gravity, or $\Gamma=0$, before adding scatters). Those extreme individuals are excluded when $\Gamma_{\rm{med}}$ is derived by consolidation in a mock sample.
Figure 25 shows the distributions of $\Gamma_{\rm{med}}$ from mock Gaia observations. The biases or shifts of $\Gamma$ are overall similar to the case of mock Newtonian observations shown in Figure 24, but show some differences in detail. Unlike the Newtonian case, the transition regime is displayed in the left column because some (though small) shifts occur. In the MOND regime, the shifts and the widths of the distributions are overall similar to those in the Newtonian case. The exception is the shift in the middle row. This appears to be due to the relatively large input value of $\Gamma$ while the median of the distribution is similar to the bottom row. In the transition+MOND regime, the widths are significantly narrower than in the Newtonian case, but the widths from the mock Newtonian and mock Gaia cases bracket the widths of the consolidated distributions shown in the right column of Figure 22. In line with Figure 22 and Figure 24, Figure 25 indicates that Newtonian gravity is inconsistent with the Gaia data, if any, even more strongly than in the other figures.
Figure 26: Each row is same as the upper row of Figure 22 but with a correction/variation (or corrections/variations) to the nominal inputs as indicated. See the text for the details.
So far all results are based on the nominal/standard inputs. Here, reasonable corrections or variations to the inputs are considered to estimate known (but ignored) systematics. They include the H2022 prior on eccentricities, perspective effects in observed velocity components, and Gaia FLAME masses. Note that perspective effects are calculated (see Appendix B) assuming that two stars in a pair are at the same distance, and thus I refer to them as first-order estimates. When the tiny distance difference is not measurable, it appears most reliable to assume that it is zero (see Appendix B for further). Note also that FLAME masses are available for about two thirds of stars, and thus nominal masses are replaced only for those stars.
Figure 26 shows results on consolidated gravity in the MOND and transition regimes with the alternative inputs. Notice that I also consider the case with both the H2022 prior and the perspective effects. The H2022 prior slightly increases $\Gamma$ compared with the thermal prior while correcting velocities for the perspective effects slightly decreases it. Thus, if both are considered, $\Gamma$ is essentially unchanged from the nominal results. FLAME masses do not lead to any tangible change either, as expected from Figure 13. In summary, any ignored known systematics lead to negligibly small variations of $\Gamma$ compared with its statistical uncertainties based on the present data.
@@ 4.1 Meaning of the results and an exploration of possible systematic errors
When the center-of-mass of a binary system orbits in the Galaxy, the relative motion between the two stars will suffer from perturbations due to long-range chance encounters with other stars and the Galactic tidal field (see Jiang & Tremaine 2010). Such perturbations can accumulate over a long period of time since the birth of the binary system (e.g. several giga years), leading to an evolved statistical property of binary populations as studied by Jiang & Tremaine (2010) in Newtonian gravity. The maximum projected separation for the wide binaries in the nominal sample (Table 1) is $\approx 25$ kau. This corresponds to $\approx 0.074r_{J}$ where $r_{J}\approx 1.65$ pc is the Jacobi radius for the median total mass of $1.82M_{\odot}$ based on Equation (43) of Jiang & Tremaine (2010). According to Newtonian simulation results shown in Figure 7 of Jiang & Tremaine (2010), presently observed binaries with $s/r_{J}\lesssim 0.074$ would be expected to follow Keplerian behavior well. This indicates that the assumption of elliptical orbits is valid as long as Newtonian gravity is tested with the wide binaries of this work. Thus, the disagreement (Figure 22) with Newton in low-acceleration regimes (along with the agreement in a strong-acceleration regime) is an unacceptable result in Newtonian gravity. Note also that the local dark matter density ($\lesssim 0.01M_{\odot}\text{ pc}^{-3}$: see Read (2014)) inferred assuming standard gravity cannot make any contribution to the gravitational anomaly.
@@
However, in MOND or other modified gravity it is unclear how good the two assumptions will be and how to interpret the Bayesian inferred values of $\Gamma$ ($>0$) (i.e.  $G_{\rm{eff}}>G_{\rm{N}}$). In MOND gravity (e.g. Bekenstein & Milgrom 1984; Milgrom 2010), the rather strong external field $\approx 1.9a_{0}$ of the Galaxy can warp the orbit and the orbit may not be closed when the external field is inclined with respect to the rotation axis (Brada & Milgrom, 2000). In such gravity models, numerical simulations are needed to produce mock wide binaries and to infer $\Gamma$ in the instantaneous motions of the binaries. The values of $\Gamma$ can then be compared with those for the Gaia binaries.
@@
Therefore, a specific modified gravity model cannot be accurately tested with the measured value of $\Gamma$. The agreement of the Bayesian-inferred values of $\Gamma=0.085\pm 0.040$ (MOND regime) and $\Gamma=0.063\pm 0.015$ (transition+MOND regime) with the generic prediction of MOND gravity ($\approx 0.07$) needs to be taken with a grain of salt. In this respect, it would be interesting to reconsider the wide binaries with wildly large values of $\Gamma$ (Table 2). These systems were rejected as systems with suspicious kinematic contaminants or abnormal RV data because they appeared to be incompatible not only with Newtonian gravity but also with MOND gravity. However, it is unclear whether MOND gravity could accommodate these systems until realistic simulations are used to probe all possible orbital motions.
Whatever the case with MOND gravity, if at least some of these systems are pure binaries without kinematic contaminants or abnormal RV data, standard gravity will be ruled out more severely. Perhaps, it is unlikely that all of them have kinematic contaminants or abnormal RV data for several reasons. (1) All of them already satisfy the presently highest data qualities and 13 of the 16 stars have ruwe $<1.2$. (2) Gaia DR3 astrometric solutions are based on data collected over a time span of 34 months, and thus any hidden close binaries with periods of the same order or shorter have already been flagged and precluded. (3) The availability of the Gaia-measured RVs with relatively good precision indicates that the stars are likely to satisfy relatively good single-star solutions for spectroscopic observations. (4) The posterior PDFs of all parameters are well-behaved and the fitted values of $e$ are particularly reasonable. Therefore, it may well be that at least some of these binaries individually rule out standard gravity.
@@ 4.2 Wide binary results in the context of recent small-scale gravity tests
Unlike the above two studies (Desmond et al., 2024; Vokrouhlický et al., 2024) on the dynamics of solar system bodies, Brown & Mathur (2023) argue for MOND gravity. Kuiper Belt objects (KBOs) with semi-major axes greater than 250 au exhibit orbital anomalies that require an unidentified planet ($5-10$ times the mass of Earth) at an average distance of about 500 au in standard gravity, but may be alternatively explained by the quadrupole and octupole moments of MOND modified gravity (Brown & Mathur, 2023) (but see Vokrouhlický et al. (2024) for possible issues with their results). Future studies of KBOs may be interesting. Also, it is interesting to note another recent small-scale test with the dynamics of open clusters (although it is not a study of the solar system). Kroupa et al. (2024) find that asymmetrical tidal tails of open clusters strongly imply gravitational anomalies that agree with MOND gravity but strongly rule out standard gravity.
At the heart of the debate between standard gravity (with dark matter) and MOND or modified gravity (without dark matter) in the universe is whether a direct gravitational boost occurs at low acceleration $\lesssim a_{0}$. The low-acceleration gravitational boost is the fundamental nature that is shared by all models belonging to the nonstandard paradigm, while solar system tests discussed above are more about specifics of models. This means that even if solar system tests truly rule out certain modified gravity models, they cannot rule out the MOND or modified gravity paradigm as long as the low-acceleration gravitational anomaly is experimentally true. This fundamental nature cannot be so easily tested with galactic dynamics (although feasible in principle and will be done eventually) because the prediction of modified gravity needs to be distinguished from that of the cosmologically inferred dark matter density (Chae, 2022). Unlike galactic dynamics, wide binary internal dynamics does not care whether dark matter density is zero or what general relativity infers cosmologically, and thus most directly tests the low-acceleration gravitational anomaly. What the results from this work and several recent studies (Chae, 2023a, 2024a, 2024c; Hernandez, 2023; Hernandez et al., 2024a; Hernandez & Kroupa, 2025) show clearly is that the fundamental low-acceleration gravitational boost is present in the currently available data.